OpenStudy (anonymous):

Solve the equation or formula for the indicated variable. S=6r^5t, for t

OpenStudy (nikato):

Divide by sides by 6r^5

OpenStudy (anonymous):

Im sorry the 5 is actually 3

OpenStudy (nikato):

*both

OpenStudy (anonymous):

so divide both sides by 3

OpenStudy (nikato):

U mean 6r^3

OpenStudy (anonymous):

yes

OpenStudy (nikato):

Yea, that's it

OpenStudy (anonymous):

\[\frac{ S }{ r^3 }=t\] ? idk if i did that right

OpenStudy (nikato):

Yea, that right. I think u forgot the 6 in ur denominator

OpenStudy (anonymous):

so, \[\frac{ S }{ 6r^3 }=t\]

OpenStudy (nikato):

Yea, that's it

OpenStudy (anonymous):

Thank you!

OpenStudy (nikato):

No problem